Identify whether the parabola y=-(x-3)^2 +6 opens up or down and identify the vertex, x-intercepts and y-intercepts. I know it opens down but, I am not sure how to identify the intercepts or vertex (3,6)?
Answer by stanbon(75887)

y=-(x-3)^2 +6
Intercepts:
Let x=0 to get the y-intercept:
y=-9+6=-3; y-intercept is (0.3)
Let y=0 to get the x-intercepts:
0=-(x-3)^2+6
(x-3)^2=6
x-3=+-sqrt6
x=3+sqrt6 or x=3-sqrt6
These are the x intercepts.
-----------------
Vertex:
(x-3)^2=-(y-6)
Vertex at (3,6)
Cheers,
Stan H.